Please help me to solve this problem . the sum of the ages of two children is 16 years . four years ago the age of the older child was three times the age of the younger child. find the present age of each child
Answer by JulietG(1812)

A + B = 16
(A-4) = 3(B-4)
Distribute
(A-4) = 3B-12
Add 4 to each side
A = 3B - 8
Substitute the value of A that you just calculated into the top equation.
A + B = 16
(3B-8) + B = 16
4B-8 = 16
Add 8 to each side
4B = 24
Divide each side by 4
B = 6
If the younger child is 6, then the older is 10 (16-6)

Prove it:
four years ago the age of the older child (10-4=6) was three times the age of the younger child (6-4=2)
6 = 3x 2